# CARE: Contrastive Alignment for ADL Recognition from Event-Triggered Sensor Streams

### v2

@@ III-C Sequence–Image Contrastive Alignment (SICA)
Our goal is to unify the complementary strengths of the sequence- and image-domain streams by learning a shared latent space. The sequence-domain stream captures temporal dependencies explicitly. The image-domain stream emphasizes spatial relationships and frequency-sensitive patterns and also preserves implicit temporal order via event sequencing. These two perspectives are complementary and partially overlapping: one models fine-grained temporal dynamics, the other grounds activities in spatial context with implicit temporal correlations.
@@
To address these limitations, we propose Sequence–Image Contrastive Alignment (SICA), which introduces a contrastive objective (i.e., the SICA loss) to explicitly align sequence and image embeddings in one shared latent space. By pulling together cross-view embeddings of the same activity and pushing apart those of different activities, SICA enforces both cross-view alignment and class-level discriminability, enabling the model to fully exploit temporal and spatially grounded cues.
Projection Heads. To enable contrastive alignment across heterogeneous representation views, we introduce lightweight projection heads after both the sequence and image encoders. Each projection head is a single-layer MLP that maps view-specific encoder outputs into a shared latent space with matched dimensionality. This transformation ensures comparability across views and refines representations for contrastive learning.
@@
Contrastive Objective. Let $P_{S}(i)$ and $P_{I}(i)$ denote the intra-view and cross-view positive sets for anchor $i$, and let $A(i)$ denote the set of all non-anchor embeddings in the batch. The instance-level loss is defined as:
$\mathcal{L}_{\mathrm{SICA},i}$ $=\frac{-1}{|P_{S}(i)|+|P_{I}(i)|}\Bigg(\sum_{p\in P_{S}(i)}\log\frac{\exp(\langle\mathbf{z}_{i},\mathbf{z}_{p}\rangle/\tau)}{\sum_{a\in A(i)}\exp(\langle\mathbf{z}_{i},\mathbf{z}_{a}\rangle/\tau)}$ (1)
$\hskip 18.49988pt\qquad+\sum_{p\in P_{I}(i)}\log\frac{\exp(\langle\mathbf{z}_{i},\mathbf{z}_{p}\rangle/\tau)}{\sum_{a\in A(i)}\exp(\langle\mathbf{z}_{i},\mathbf{z}_{a}\rangle/\tau)}\Bigg),$
@@ III-D CARE: An End-to-End Framework of Contrastive Alignment for ADL Recognition
- Contrastive Representation Learning Branch: This branch implements SICA. Encoders ($Enc_{S}(\cdot),Enc_{I}(\cdot)$) produce view-specific features ($\mathbf{r}_{i}^{S},\mathbf{r}_{i}^{I}$), which are projected into a shared latent space via lightweight projections ($Proj_{S}(\cdot),Proj_{I}(\cdot)$). The SICA loss $\mathcal{L}_{SICA}$ enforces intra-view consistency and cross-view alignment. It is worth to note that both the sequence encoder and image encoder in CARE are modular components that can be flexibly replaced with alternative encoder architectures.
@@
where $0\leq\beta\leq 1$ balances cross-view alignment and supervised classification. The classification head not only enables direct prediction but also provides gradients that complement SICA, improving representation discriminability in an end-to-end manner.
@@ IV-A Datasets and Experimental Protocol
Datasets. We evaluate on three CASAS datasets—Milan, Cairo, and Kyoto7—collected in realistic smart-home testbeds with bedrooms, bathrooms, kitchens, living rooms, and dining rooms instrumented by ambient sensors (motion, temperature, door), following Cook et al. [13]. Milan contains 35 sensors, 15 activity categories, and 4,060 recorded activities; Cairo has 34 sensors, 13 categories, and 842 activities; Kyoto7 includes 71 sensors, 13 categories, and 634 activities. For comparability, we group the original labels into 11 ADLs, consistent with [11].
Data splits. We adopt a stratified 70/30 train/test split by activity class. On the 70% training portion, we run 5-fold cross-validation for hyperparameter tuning, then retrain on the full training split and evaluate on the held-out test split. Each experiment is repeated with 5 random seeds; we report mean $\pm$ std.
Batching and training. Batch size is 64; training runs up to 60 epochs with early stopping on validation F1. No random window sampling is used.
Encoders. The sequence/image encoders in Sec. III-A and Sec. 3 are architecture-agnostic and can be instantiated with any sequence or image backbone. In our experiments, we instantiate them with LSTM or BiLSTM for the sequence stream and ResNet18 for the image stream; the corresponding dual-stream variants are denoted L-Res and BiL-Res.
@@ IV-B Overall Performance
Baselines. We compare our method against five categories of baselines: (1) Traditional ML: Random Forest, SVM; (2) Sequence-based: DeepCASAS [26], Embedding FCN [4], TCN [22], GRU [10], 1D-CNN [25], LSTM-CNN [37]; (3) Image-based: DCNN [18], GreyDCNN [28], Transformer [14]; (4) Graph-based: GraphModel [31]; (5) Language-inspired: TDOST [34], which frames activity recognition under a language modeling paradigm. This diverse set of baselines covers traditional feature-based classifiers, sequence/image/graph encoders, and recent language-inspired models, providing a comprehensive benchmark for evaluating our framework. All baseline and proposed models are trained and evaluated under the exact same train/test splits and evaluation protocol, ensuring fairness and eliminating confounding factors due to differing data partitions.
Overall performance. Table II compares all approaches on Milan, Cairo, and Kyoto7. Traditional machine learning models, such as Random Forest and SVM, perform significantly worse than deep learning baselines, underscoring the necessity of representation learning in complex ADL dynamics. Deep learning approaches yield substantially higher performance. Sequence-based baselines (e.g., DeepCASAS, EmbeddingFCN, GRU, LSTM-CNN) achieve competitive results as they usually capture temporal dependencies effectively, but are sometimes more affected by noise and sparsity. While image-based baselines (e.g., BinaryCNN, DCNN, GreyDCNN, Transformer) are generally more robust but may degrade on highly sparse datasets such as Kyoto7. This complementarity highlights the importance of integrating both views. The graph-based model performs competitively on Milan and Cairo by explicitly modeling sensor relationships, but suffers a marked performance degradation on the sparse and noisy kyoto7. By using a textual encoding method, the accuracy of TDOST is generally better than that of all the other baselines except ours, demonstrating the advantage of textual information in capturing cross-view information. Overall, our SICA models consistently deliver superior performance across all datasets and metrics, establishing a new state of the art.
Computation considerations. In general, image-based models require more computation than sequence-based ones due to the inherent differences in input size and model complexity. GraphModel further exacerbates this with prohibitively high inference time (1.586 s/batch). TDOST consumes only 0.001 seconds per batch by using BiLSTM to inference. Despite leveraging both representation views, our SICA models remain efficient (0.010–0.011 s/batch), on par with image-based baselines and an order of magnitude faster than GraphModel. This favorable trade-off between accuracy and efficiency underscores the scalability of SICA for real-world smart-home ADL recognition.
@@
Metric Uni-View Within-View Align Cross-View Align
LSTM BiLSTM ResNet18 L-Res BiL-Res L-Res BiL-Res
Milan
@@
Eff.
Mem. 171.76 313.53 1511.19 1620.17 1716.93 1620.17 1716.93
T/B 0.004 0.005 0.006 0.010 0.011 0.010 0.011
We evaluate three variants: (i) uni-view baselines (LSTM, BiLSTM, ResNet18), which learn within a single domain; (ii) within-view alignment, which applies contrastive loss only to intra-view pairs; and (iii) cross-view alignment (our full SICA), which leverages both intra- and cross-view pairs.
As shown in Tab. III, cross-view alignment consistently outperforms all baselines. For instance, BiL-Res achieves gains of $\sim$3% on Milan, 2% on Cairo, and 1% on Kyoto7 compared with the best unimodal models. Compared to within-view alignment, cross-view alignment further improves accuracy by 1%, confirming the value of explicitly enforcing consistency across heterogeneous views.
@@
Fig. 6 compares the two approaches across three datasets. SICA have a comparable performance with CE-based fusion, with the largest gains observed on the challenging Kyoto7 dataset. For example, BiL-Res with SICA improves the F1-score by 1.3% over its CE counterpart. Importantly, these CE models are not previous baselines but architectural counterparts of our framework, differing only in the training objective—underscoring the benefit of contrastive alignment.
@@ IV-E Temporal Binning and Frequency-based Event Filtering
Temporal binning. We first examine the impact of discretizing timestamps into coarse-grained bins. Fig. 7 reports results with bin widths of 0.5h, 1h, and 4h ($\Delta T\in\{0.5,1,4\}$), compared to the case without incorporating time information. The results without incorporating time information is lower than all the other settings by 1–3%, suggesting the importance of timestamps. A 1-hour bin consistently yields the best trade-off across all datasets: finer bins (0.5h) introduce redundancy and unstable variance (e.g., on Kyoto7), while coarser bins (4h) discard critical context between adjacent activities. This aligns with the average activity durations (0.25h for Milan, 1.09h for Cairo, and 0.67h for Kyoto7), confirming that the 1-hour setting best balances temporal resolution and robustness.
@@
Consequently, FEF eliminates sensors that are highly irrelevant to the target activity; however, it may also remove useful information. Therefore, an appropriate filtering threshold must be carefully chosen. We also test the effect of removing sensors with very low activation frequency. As shown in Table IV, the best performance on Cairo is achieved with a 1% threshold; results obtained without FEF or at higher thresholds exhibit a slight decline. This suggests that filtering suppresses spurious noise while preserving salient sensor events, with 1% threshold providing the best robustness–information balance.
@@
TABLE IV: BiL-Res on Cairo: Frequency-based Event Filtering (FEF) Threshold $\theta$.
@@
Together, these ablations confirm that our preprocessing pipeline—hourly temporal binning plus Frequency-based Event Filtering—strikes the right compromise between noise suppression and information retention, providing a clean foundation for downstream multimodal representation learning.
@@ IV-F Effectiveness of Image-Based Representations
Figure 10: Cairo’s Grad-Cam on Work Activity.
IV-G Robustness to Sensor Corruption and Layout Variability
In real-world deployments, sensor malfunctions, missing signals, and changes in sensor placement are inevitable sources of distribution shift. To evaluate robustness under such conditions, we test our models trained on clean data directly on corrupted test sets without retraining.
@@ IV-G Robustness to Sensor Corruption and Layout Variability
Results. As shown in Tab. V, our BiL–Res model exhibits strong resilience: accuracy on Milan, Cairo and Kyoto7 drops by only 1–3% across all corruption settings.

### v3

@@ III-C Sequence–Image Contrastive Alignment (SICA)
Our goal is to unify the complementary strengths of the sequence- and image-domain streams by learning a shared latent space. The sequence-domain stream captures temporal dependencies explicitly. The image-domain stream emphasizes spatial relationships and frequency-sensitive patterns, and preserves implicit temporal order via event sequencing. These two perspectives are complementary and partially overlapping: one models fine-grained temporal dynamics, the other grounds activities in spatial context with implicit temporal correlations.
@@
To address these limitations, we propose Sequence–Image Contrastive Alignment (SICA), which introduces a contrastive objective (i.e., the SICA loss) to explicitly align sequence and image embeddings in a shared latent space. By pulling together cross-view embeddings of the same activity and pushing apart those of different activities, SICA enforces cross-view alignment and class-level discriminability, enabling the model to exploit temporal and spatially grounded cues.
Projection Heads. To enable contrastive alignment across heterogeneous representation views, we introduce lightweight projection heads after sequence and image encoders. Each projection head is a single-layer MLP that maps view-specific encoder outputs into a shared latent space with matched dimensionality. This transformation ensures comparability across views and refines representations for contrastive learning.
@@
Contrastive Objective. Let $P_{in}(i)$ and $P_{cr}(i)$ denote the intra-view and cross-view positive sets for anchor $i$, and let $A(i)$ denote the set of all non-anchor embeddings in the batch. The instance-level loss is defined as:
$\mathcal{L}_{\mathrm{SICA},i}$ $=\frac{-1}{|P_{in}(i)|+|P_{cr}(i)|}\Bigg(\sum_{p\in P_{in}(i)}\!\!\log\frac{\exp(\langle\mathbf{z}_{i},\mathbf{z}_{p}\rangle/\tau)}{\sum_{a\in A(i)}\exp(\langle\mathbf{z}_{i},\mathbf{z}_{a}\rangle/\tau)}$ (1)
$\hskip 18.49988pt\qquad+\sum_{p\in P_{cr}(i)}\log\frac{\exp(\langle\mathbf{z}_{i},\mathbf{z}_{p}\rangle/\tau)}{\sum_{a\in A(i)}\exp(\langle\mathbf{z}_{i},\mathbf{z}_{a}\rangle/\tau)}\Bigg),$
@@ III-D CARE: An End-to-End Framework of Contrastive Alignment for ADL Recognition
- Contrastive Representation Learning Branch: This branch implements SICA. Encoders ($Enc_{S}(\cdot),Enc_{I}(\cdot)$) produce view-specific features ($\mathbf{r}_{i}^{S},\mathbf{r}_{i}^{I}$), which are projected into a shared latent space via lightweight projections ($Proj_{S}(\cdot),Proj_{I}(\cdot)$). The SICA loss $\mathcal{L}_{SICA}$ enforces intra-view consistency and cross-view alignment. It is worth noting that both the sequence encoder and image encoder in CARE are modular components that can be flexibly replaced with alternative encoder architectures.
@@
where $0\!\leq\!\beta\!\leq\!1$ balances cross-view alignment and supervised classification. The classification head enables direct prediction and also provides gradients that complement SICA, improving representation discriminability in an end-to-end manner.
@@ IV-A Datasets and Experimental Protocol
Datasets. We evaluate on three CASAS datasets—Milan, Cairo, and Kyoto7—collected in realistic smart-home testbeds with bedrooms, bathrooms, kitchens, living rooms, and dining rooms instrumented by ambient sensors (motion, temperature, door), following Cook et al. [9]. Milan contains 35 sensors, 15 activity categories, and 4,060 recorded activities; Cairo has 34 sensors, 13 categories, and 842 activities; Kyoto7 includes 71 sensors, 13 categories, and 634 activities. For comparability, we group the original labels into 10 ADLs with an additional null (“Other”) class, following the standard protocol in prior work (e.g., [24, 32, 16, 26, 29]).
Data splits. We adopt stratified 70/30 train/test split by activity class. On training portion, we run 5-fold cross-validation for hyperparameter tuning, then retrain on the full training split and evaluate on the held-out test split. Each experiment is repeated with 5 random seeds; we report mean $\pm$ std.
Batching and training. Batch size is 64; training runs up to 60 epochs with early stopping on validation F1. No random window sampling is used. We use a learning rate of 0.001, weight decay of 0.001, and no learning-rate scheduling. The temperature of the supervised contrastive loss is set to 0.05.
Model structure. Our sequence/image encoders (Sec. III-A, Sec. III-B) are architecture-agnostic that can be instantiated with any sequence or image backbone. In experiments, we use LSTM or BiLSTM for the sequence stream and ResNet18 for the image stream; the corresponding dual-stream variants are denoted L-Res and BiL-Res. Both streams are projected to a 32-d shared embedding for contrastive alignment, and an MLP classifier predicts labels from the concatenated features. Implementation details are in the released codebase.
@@ IV-B Overall Performance
Baselines. We compare our method against five categories of baselines: (1) Traditional ML: Random Forest, Support Vector Machines (SVM); (2) Sequence-based: DeepCASAS [24], Embedding FCN [4], TCN [23], GRU [8], 1D-CNN [22], LSTM-CNN [35]; (3) Image-based: DCNN [16], GreyDCNN [26], Transformer [10]; (4) Graph-based: GraphModel [29]; (5) Contrastive-based: CPC [18]; (6) Language-inspired: TDOST [32], which frames activity recognition under a language modeling paradigm. This diverse set of baselines covers traditional feature-based classifiers, sequence/image/graph encoders, and recent language-inspired models, providing a comprehensive benchmark for evaluating our framework. All baseline and proposed models are trained and evaluated under the exact same train/test splits and evaluation protocol, ensuring fairness and eliminating confounding factors due to differing data partitions.
Overall performance. Table II compares all approaches on Milan, Cairo, and Kyoto7. Traditional machine learning models, such as Random Forest and SVM, perform significantly worse than deep learning baselines, underscoring the necessity of representation learning in complex ADL dynamics. Deep learning approaches yield substantially higher performance. Sequence-based baselines (e.g., DeepCASAS, EmbeddingFCN, GRU, LSTM-CNN) achieve competitive results as they usually capture temporal dependencies effectively, but are sometimes more affected by noise and sparsity. While image-based baselines (e.g., BinaryCNN, DCNN, GreyDCNN, Transformer) are generally more robust but may degrade on highly sparse datasets such as Kyoto7. This complementarity highlights the importance of integrating both views. The graph-based model performs competitively on Milan and Cairo by explicitly modeling sensor relationships, but suffers a marked performance degradation on the sparse and noisy Kyoto7. By using a textual encoding method, the accuracy of TDOST is generally better than that of all the other baselines except ours, demonstrating the advantage of textual information in capturing cross-view information. The contrastive-based method, CPC, shows competitive performance on Kyoto7, but falls short on Milan and Cairo compared with our method and TDOST. Overall, our SICA models consistently deliver superior performance across all datasets and metrics, establishing a new state of the art.
Computation considerations. In general, image-based models require more computation than sequence-based ones due to the inherent differences in input size and model complexity. GraphModel further exacerbates this with prohibitively high inference time (1.586 s/batch). TDOST consumes only 0.001 seconds per batch by using BiLSTM for inference. Despite leveraging both representation views, our SICA models remain efficient (0.010–0.011 s/batch), on par with image-based baselines and an order of magnitude faster than GraphModel. This favorable trade-off between accuracy and efficiency underscores SICA’s scalability for real-world ADL recognition.
@@
Metric Uni-View Within-View Align Self-Supervised Align Cross-View Align
LSTM BiLSTM ResNet18 L-Res BiL-Res L-Res BiL-Res L-Res BiL-Res
@@
Eff. Mem. 171.76 313.53 1511.19 1620.17 1716.93 1620.17 1716.93 1620.17 1716.93
T/B 0.004 0.005 0.006 0.010 0.011 0.010 0.011 0.010 0.011
We evaluate four variants: (i) uni-view baselines (LSTM, BiLSTM, ResNet18), which learn within a single domain; (ii) within-view alignment, which applies contrastive loss only to intra-view pairs; (iii) cross-view alignment (our full SICA), which leverages both intra- and cross-view pairs; and (iv) self-supervised alignment can be regarded as an instance-level cross-view alignment, which aligns the two views using an instance-level contrastive objective: only the paired embeddings of the same sample are treated as positive pairs, without using class labels to construct cross-sample positives.
As shown in Tab. III, cross-view alignment consistently outperforms all baselines. For instance, BiL-Res achieves gains of $\sim$3% on Milan, 2% on Cairo, and 1% on Kyoto7 compared with the best unimodal models. Compared with within-view alignment, cross-view alignment yields comparable or up to $\sim 1\%$ higher accuracy, and is our default due to its consistent gain across datasets. Instance-level cross-view alignment is competitive on Milan but degrades on smaller datasets. Both validate the effectiveness of our cross-view design.
@@
Fig. 6 compares the two approaches across three datasets. SICA has a comparable performance with CE-based fusion, with the largest gains observed on the challenging Kyoto7 dataset. For example, BiL-Res with SICA improves the F1-score by 1.3% over its CE counterpart. Importantly, these CE models are not previous baselines but architectural counterparts of our framework, differing only in the training objective—underscoring the benefit of contrastive alignment.
@@
Furthermore, in Eq. III-D, the hyperparameter $\beta$ controls the extent to which SICA deviates from naive fusion. We evaluate BiL-Res under different $\beta$ and report the corresponding mean $\pm$ std bounds in Fig. 7. Across all datasets, performance varies within approximately 1% on Milan, 2% on Cairo, and 4% on Kyoto7, demonstrating that CARE is relatively insensitive to the choice of $\beta$ and robust to moderate variations of it.
Figure 7: Sensitivity Analysis of $\beta$ for BiL-Res.
@@ IV-E Temporal Binning and Frequency-based Event Filtering
Temporal binning. We first examine the impact of discretizing timestamps into coarse-grained bins. Fig. 8 reports results with bin widths of 0.5h, 1h, and 4h ($\Delta T\in\{0.5,1,4\}$), compared to the case without incorporating time information. The results without incorporating time information is lower than all the other settings by 1–3%, suggesting the importance of timestamps. A 1-hour bin consistently yields the best trade-off across all datasets: finer bins (0.5h) introduce redundancy and unstable variance (e.g., on Kyoto7), while coarser bins (4h) discard critical context between adjacent activities. This aligns with the average activity durations (0.25h for Milan, 1.09h for Cairo, and 0.67h for Kyoto7), confirming the 1-hour setting best balances temporal resolution and robustness. We also find that without temporal binning, the model remains competitive, suggesting binning is not the sole driver of performance.
@@
Consequently, FEF eliminates sensors that are highly irrelevant to the target activity; however, it may also remove useful information. Therefore, an appropriate filtering threshold must be carefully chosen. We also test the effect of removing sensors with very low activation frequency. As shown in Table IV, the best performance on Cairo is achieved with a 1% threshold; results obtained without FEF or at higher thresholds exhibit a slight decline. This suggests that filtering suppresses spurious noise while preserving salient sensor events, with 1% threshold providing the best robustness–information balance. To further isolate FEF’s contribution, we integrate it into DeepCASAS. This boosts accuracy by $\sim$3% on Milan and $\sim$8% on Kyoto7, supporting FEF’s effectiveness and generality. In contrast, the gain on Cairo is marginal (<<1%), suggesting preprocessing alone cannot account for the improvement. Collectively, the gains are driven by CARE’s overall design.
TABLE IV: BiL-Res on Cairo: FEF Threshold $\theta$.
@@
We acknowledge that the temporal bin width and frequency thresholds are manually selected, which may limit the generality of the method. However, ablations show the method is not sensitive to their choices (accuracy varies by $3\%$ across datasets), so coarse tuning is sufficient in practice). Together, the ablations confirm that our preprocessing pipeline—hourly temporal binning plus Frequency-based Event Filtering—strikes the right compromise between noise suppression and information retention, providing a clean foundation for downstream multimodal representation learning.
@@ IV-F Effectiveness of Image-Based Representations
IV-G Robustness to Sensor Corruption, Layout Variability and Temporal Drift
Real-world deployments face distribution shifts from sensor malfunctions, missing signals, sensor layout changes, and temporal drift. We train on clean data and evaluate directly on corrupted test sets without retraining to test robustness to corruption and layout variability. For temporal drift, we adopt a time-aware split: train on the earlier $70\%$ and test on the remaining $30\%$.
@@ IV-G Robustness to Sensor Corruption, Layout Variability and Temporal Drift
TABLE VI: DeepCASAS and TDOST with Corrupt($5\%\times 5\%$)
Baseline Milan Cairo Kyoto7
Accuracy F1-Score Accuracy F1-Score Accuracy F1-Score
[…]
Results. As shown in Tab. V, our BiL–Res model exhibits strong resilience: accuracy on Milan, Cairo and Kyoto7 drops by only 1–3% across all corruption settings. We further conducted robustness evaluations of DeepCASAS and TDOST under the $5\%\times 5\%$ corruption in Tab. VI. Overall, CARE is consistently robust across datasets and degrades less than strong baselines under the same corruption setting.
@@
TABLE VII: Performance Comparison under Time-aware Split
Model Milan Cairo Kyoto7
Accuracy F1-Score Accuracy F1-Score Accuracy F1-Score
DeepCASAS (LSTM) $74.7\pm 0.9$ $71.0\pm 1.0$ $63.9\pm 2.0$ $55.5\pm 3.0$ $51.6\pm 7.6$ $48.0\pm 9.2$
DeepCASAS (BiLSTM) $75.7\pm 0.5$ $72.5\pm 0.6$ $68.7\pm 4.8$ $62.8\pm 7.5$ $55.1\pm 1.2$ $51.4\pm 1.7$
LSTM-ResNet18 80.3 $\pm$ 1.3 78.2 $\pm$ 1.5 73.9 $\pm$ 2.5 72.6 $\pm$ 3.5 68.6 $\pm$ 1.9 67.1 $\pm$ 2.9
BiLSTM-ResNet18 $77.9\pm 0.4$ $75.8\pm 1.1$ $73.2\pm 3.2$ $70.7\pm 2.0$ $68.1\pm 4.5$ $65.3\pm 6.8$
Tab. VII reports results under the time-aware split. Despite temporal-drift degradation for both methods, ours consistently outperforms DeepCASAS by about $2\%$ on Milan, $5\%$ on Cairo, and $13\%$ on Kyoto7.